Circular No. 8054

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V838 MONOCEROTIS D. K. Lynch, R. W. Russell, and D. Kim, The Aerospace Corporation;

and M. L. Sitko, University of Cincinnati, report 3–14-μm spectroscopy of V838 Mon on Jan. 8.4 UT, about a year after visual maximum brightness: “Using the Aerospace Broadband Array Spectrograph System (BASS) on the NASA Infrared Telescope Facility 3-m telescope, we observed a spectrum dominated by what appear to be molecular absorption/emission bands, most of which are unidentified except for water near 3 μm. A quasicontinuum was apparent with a temperature of ∼ 1000 K (with considerable temperature uncertainty). Narrow-band magnitudes (± 0.05) were M = 2.7 and N = 1.3. This represents a significant brightening and shape change since our previous observations with the same equipment on 2002 Jan. 15.49. At that time, the spectrum was smooth and consistent with a black body at ∼ 2350 K, except between 8 and 10 μm, which showed a roughly-50-percent excess. The narrow-band magnitudes were L = 5.3, M = 5.2 and N = 5.2, each ± 0.1.”

XTE J1720–318 M. P. Rupen, National Radio Astronomy Observatory (NRAO); C.

Brocksopp, Mullard Space Science Laboratory; A. J. Mioduszewski, V. Dhawan, NRAO; R. Sault, Australia Telescope National Facility; R. Fender, University of Amsterdam; and S. Corbel, University of Paris and Laboratoire du Commissariat à l’Énergie Atomique, Saclay, report the detection of a radio counterpart of XTE J1720–318 (IAUC 8050). Observations were made with the Very Large Array (VLA) on Jan. 15 UT, and with the Australia Telescope Compact Array on Jan. 16. There are several radio sources within the x-ray error circle, but only one of these has varied significantly, rising at 4.9 GHz from 0.32 ± 0.04 mJy on Jan. 15 to 4.9 ± 0.1 mJy on Jan. 16. The VLA data give a position of α = 17^{h}19^{m}59^{s}.062 ± 0^{s}.087, δ = −31^{o}44′59′′.7 ± 1′′.10 (equinox 2000.0), ∼ 1′.4 from the original x-ray position. Further VLA data from Jan. 15 give a flux density at 8.5 GHz of 0.27 ± 0.06 mJy, and a non-detection at 1.4 GHz (rms of 0.24 mJy/beam). The Northern VLA Sky Survey (Condon et al. 1998, A.J. 115, 1693) also show no source at 1.4 GHz in 1995, to an rms noise limit of 0.52 mJy/beam. Further radio observations are planned, and observations at other wavelengths are urgently requested.

2003 January 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green